A rectangular garden is 10ft longer than its width. Its area is 875ft2. What are its dimensions?
Answer by stanbon(75887)

A rectangular garden is 10ft longer than its width. Its area is 875ft2. What are its dimensions?
----
Let width be "x"
Then length = "x+10"
----
Equation:
Area = L*W
x(x+10) = 875
x^2 + 10x -875 = 0
--------------
(x-25)(x+35) = 0
width = x = 25 ft
length = x+10 = 35 ft
=============
Cheers,
Stan H.
===============